Suppose there are 50 homeowners living on the lakeshore, and that each has MB = 25 - 5x, where x is the number of tankful that are sprayed. Suppose that the cost of the mosquito spray is $250 for every tankful and that mosquito spray is a pure public good. How much would each homeowner pay in a Lindahl equilibrium?

options

$0.50

$5.00

$50.00

$500.00
Answer #1

Marginal benefit=25-5X

P=25-5X

since there are 50 homeowner then market demand=1250-250X

at optimality, Price=MC=250=1250-250X

Thus 250X=1000. X=4 and each one pays MB=25-5(4)=5

ans is B